OpenStudy (anonymous):

Find the circumference of the taxicab circle with a radius of 9.

OpenStudy (anonymous):

circumference =2*pi*r so pi =22/7

OpenStudy (anonymous):

r= radius .can do that

OpenStudy (anonymous):

2*22/7*9=?